perform the indicated operation, using polar form.
4cis pie * 3cis(-pie over 2)

b) 2(cos pie + i sin pie) * 5( cos 3pie over 4 + i sin 3pie over 4)

Answer by stanbon(75887)

perform the indicated operation, using polar form.
4cis (pie) * 3cis(-pie over 2) = 12*cis(pi/2)
b) 2(cos pie + i sin pie) * 5( cos 3pie over 4 + i sin 3pie over 4)
= 10(cis(pi + (3/4)pi))
= 10(cis((7/4)pi)
=====================
Cheers,
Stan H.